Observers: Tania Moraga, Cristina Romero-Cañizales, Lluís Galbany
TAT: Joe Anderson, Morgan Fraser
Data reducers: Thomas De Jaeger, Francisco Förster
******* LOG *******
21:20 UT EFOSC bias
21:29 UT EFOSC arcs
Conditions at the beginning of the night:
Cirrus clouds / DIMM seeing: 0.78" / Humidity ~ 30% / Wind speed < 10 m/s
22:40 UT EFOSC skyflat UBVRi
- Wind btw 12 and 15 m/s from North: we cannot observe sources with DEC~>-30
23:23 UT ASASSN-15lh Gr#11 and Gr#16 / seeing=0.98 & wind=9.6 m/s
at ~00:10 conditions have gotten worse: seeing~1.8 & wind~12.6 m/s
(although now sky is clearer). The seeing varies rapidly between
1.3 to 1.8 (approx). Before Gr#16 starts, we change the slit to 1.5"
01:40 UT While the OB for ASASSN-15lh is finishing, the wind has
considerably increased (>20 m/s, average ~17 m/s) --> dome closes!